OpenStudy (clamin):

HELP!! 7x^2+5=530 solve by taking square roots

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large 7x^2+5=530\] \[\Large 7x^2+5-5=530-5\] \[\Large 7x^2=525\] \[\Large \frac{7x^2}{7}=\frac{525}{7}\] \[\Large x^2=75\] and I'll let you finish

OpenStudy (clamin):

the choices are:: a) {i square root 3, -i square root 3} b) {i square root 5} c) {5 square root 3, -5 square root 3} d) {5 square root 3}

OpenStudy (clamin):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you see how I got to x^2 = 75

OpenStudy (clamin):

yea,, you divide 525 and 7

jimthompson5910 (jim_thompson5910):

so we have x^2 = 75, what is x equal to? how do we find x?

OpenStudy (clamin):

by the 7x^2 divide by 7

jimthompson5910 (jim_thompson5910):

\[\Large x^2 = 75\]

jimthompson5910 (jim_thompson5910):

what undoes squaring?

OpenStudy (clamin):

i dont know

OpenStudy (clamin):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

example take the number 5 and square it to get 25 to undo that, you square root 25 to get 5 back

jimthompson5910 (jim_thompson5910):

so if we square x, we do what to undo that operation?

OpenStudy (clamin):

im just confuse cuz i got 75 but its not on the choices

jimthompson5910 (jim_thompson5910):

you take the square root of that to isolate x \[\Large x = \pm\sqrt{75}\]

jimthompson5910 (jim_thompson5910):

you simplify from there